OpenStudy (anonymous):

Identify the quadratic equation by the x coordinates of its roots. x = 4, x = -2

OpenStudy (anonymous):

x=4 (x-4)=0 is a factor x=-2 (x+2)=0 is a factor (x-4)(x+2)=0 just multiply them and you will get your answer